Cost of Equity: Dividend Growth Summerdahl Resort's common stock is currently trading at $24.00 a share. The stock is expected to pay a dividend of $2.00 a share at the end of the year (D1 = $2.00), and the dividend is expected to grow at a constant rate of 7% a year. What is the cost of common equity? Round your answer to two decimal places.?
Answer #1

Using DD Model,

Stock Price = D1/(r - g)

So,

24 = 2/(r - 0.07)

r = 2/24 + 0.07

r = 0.1533

Cost of Equity = 15.33%